Consider the reaction Cr2O72−(aq) + I(aq) → Cr3+(aq) + I2(s).

PWX4ljOhHSwAAAAASUVORK5CYII=
Answer #1

I2/I- is anode , E0 I2/I-  = 0.54 V

Anode reaction;

2I- - 2e \small \to I2 (1)

and Cr2O72-/Cr3+ is cathode, E0Cr2O72-/Cr3+ = 1.33 V

Cathode reaction;

Cr2O72- + 14H+ + 6 e \small \to 2 Cr3+ + 7H2O (2)

Eq. 1 * 3 + Eq. 2

Cr2O72- + 14H+ + 6 I-  \small \to 2 Cr3+ + 7H2O + 3 I2

number of electron transferred (n) = 6

E0cell = Eocathode - E0anode = 1.33 - 0.54 = 0.79 V

G0 = - nFE0cell

= - 6 \small \times 96500 \small \times 0.79

= - 457410 J

G0 = - RTlnK

or, - 457410 = - 8.314 \small \times 298.15 ln K

or, ln K = 184.52

or, K = 1.4 \small \times1080

cell reaction

Cr2O72- (aq)  + 14H+ (aq) + 6 I- (aq)  \small \to 2 Cr3+ (aq)  + 7H2O (l) + 3 I2 (s)

nernst equation of the reaction is

Ecell = E0cell -  0.0592 6 log Cr12 Cr2O H+] 14 -

Or, Ecell = 0.79 - 0.00986 log (4.28 x 102 X (2.26 x 10-2)(2.31 x 10-3)14 1(3.55 x 10-3)6 X X X

or, Ecell = 0.79 - 0.00986 log (3.29\small \times 1050)

or, Ecell = 0.79 - 0.00986\small \times 50.52

or, Ecell = (0.79 - 0.50) V

or, Ecell = 0.29 V